# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Inhomogeneous Coefficients

### v1

@@ Abstract
We consider the adaptive-rank integration of general time-dependent advection-diffusion partial differential equations (PDEs) with spatially variable coefficients. We employ a standard finite-difference method for spatial discretization coupled with diagonally implicit Runge-Kutta schemes for temporal discretization. The fully discretized scheme can be written as a generalized Sylvester equation, which we solve with an adaptive-rank algorithm structured around three key strategies: (i) constructing dimension-wise subspaces based on an extended Krylov strategy, (ii) developing an effective Averaged-Coefficient Sylvester (ACS) preconditioner to invert the reduced system for the coefficient matrix efficiently, and (iii) efficiently computing the residual of the equation without explicitly reverting to the full-rank form. The proposed approach features a computational complexity of $\mathcal{O}(N{r^{2}}+{r^{3}})$, where $r$ represents the rank during the Krylov iteration and $N$ is the resolution in one dimension, commensurate with the constant-coefficient case [10]. We present numerical examples that illustrate the computational efficacy and complexity of our algorithm.
@@
Keywords: Adaptive-rank, extended Krylov subspaces, implicit Runge-Kutta integrators, advection-diffusion equations, Sylvester preconditioning, matrix-equations.
1 Introduction
We propose an adaptive-rank numerical scheme for approximating the solution of the time-dependent variable-coefficient advection-diffusion equation. A challenge in existing numerical schemes is the curse of dimensionality, where the problem dimension $d$ leads to exponential growth in computational complexity.
These formulations often result in large matrix systems of the form $Ax=b$, with $A$ sparse and banded, of size $N^{d}$-by-$N^{d}$ and bandwidth $\mathcal{O}(N^{d-1}$), where $N$ is the resolution in one dimension and $d$ is the dimensionality. Direct solvers for such systems scale in general as $\mathcal{O}(N^{3d-2})$ [3]. The complexity of unpreconditioned Conjugate Gradient (a type of Krylov method) for diffusion problems scales as $\mathcal{O}(N^{3d/2})$ [6, 19]. Multigrid-preconditioned Krylov methods may improve this scaling to $\mathcal{O}(N^{d})$ [16], which is considered optimal. In all these cases, the presence of the dimension $d$ in the exponent makes computation of the inverse matrix challenging for fine-mesh resolutions of high-dimensional problems.
Our method seeks to overcome this dimensional scaling and provide an efficient approach with a complexity of $\mathcal{O}(N)$ to solving time-dependent variable coefficient advection-diffusion partial differential equations (PDEs). In two dimensions, our approach leverages adaptive-rank approximations of matrix solutions on a tensor-product grid. Adaptive-rank techniques have proven effective in reducing the computational complexity of high-dimensional problems by representing the solution matrix as a sum of the outer products of basis vectors. Specifically, we employ a truncated singular value decomposition (SVD), such that $X=USV^{\top}$, where basis vectors in $U$ and $V$ span the solution space and determine the rank of the solution. The key idea is to dynamically update the basis vectors, as well as the rank (denoted as $r$), to maintain accuracy while minimizing computational cost.
@@ 1 Introduction
Two primary strategies in the literature exploit low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [17, 22, 5, 8, 21] and the step-and-truncate (SAT) method, which has both explicit [18, 7, 13, 12, 14] and implicit [29, 24, 10, 23] formulations. Our work aligns with the SAT framework, where we perform temporal discretization with high-order accuracy and fully implicitly, leading to matrix differential equations to be solved at every time step. For constant coefficients, this system becomes a Sylvester equation of the form:
@@
where $\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represents the solution at time $t_{1}$, $\mathbf{X}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$ is the initial condition at $t_{0}$, and $A$ and $B$ are difference operators arising from the PDE discretization.
Krylov subspace methods have proven effective for solving Sylvester-type equations resulting from the discretization of PDEs [20, 33, 29, 25, 2]. In these methods, the solution bases are chosen from a Krylov (or Krylov-like) subspace of $A$ and $B$ in (1), and the original system is projected onto this reduced subspace using the Galerkin projection. To accelerate the convergence of iterative methods, various extensions of Krylov subspaces have been developed. Notably, the extended Krylov subspace [9] involving $A^{-1}$, $B^{-1}$ has been applied to matrix equations in [33] and later adopted for time-dependent PDEs in [10].
While substantial progress has been made in constant-coefficient problems, systems with variable coefficients remain challenging. Discretizing such systems leads to the so-called generalized Sylvester matrix equation, of the form:
@@
where $A_{i}$ and $B_{i}$ represent discretization operators for advection and/or diffusion terms. While these types of matrix equations can be solved by reverting them to the Kronecker form, in [26] the authors argue that, for a chosen approximation space, the Galerkin projection applied to the matrix equation consistently achieves faster convergence than the Kronecker formulation, particularly when the operators are poorly conditioned. Both analytical and numerical evidence in [26] support this claim. Furthermore, they show that the optimality of the Galerkin projection technique for Sylvester equations extends naturally to the generalized Sylvester equations. That is, regardless of the approximation space, the residual decreases as the subspaces are iteratively expanded, though the convergence rate depends significantly on the choice of subspaces, as demonstrated in [10]. Additional validation of this claim is provided in [34], where a matrix-based conjugate gradient method is compared with Galerkin projection techniques. Such results encourage the community to adopt matrix-equation-based strategies for multi-term matrix equations, rather than defaulting to Kronecker formulations.
@@
Several methods have been proposed to solve generalized Sylvester equations iteratively directly in matrix form, either with stationary iterative techniques or Krylov methods. Stationary iterative methods [1, 32] leverage low-rank techniques to enhance efficiency, achieving computational complexity as low as $\mathcal{O}(N)$ when implemented effectively [32]. However, their effectiveness strongly depends on whether or not such a dominating operator could be correctly identified to enable a contractive iterative process. The difficulty of explicitly finding such a dominating term in a general setting was discussed in [25], which proposed using the Kronecker formulation instead.
Krylov methods have also been explored as an alternative for the solution of equations of the Sylvester and generalized Sylvester type. In [1], the authors present a series of techniques to tackle the Generalized Lyapunov equation, which is a special case of (2) where the first two terms are of Sylvester type, i.e. $A_{1}=B_{2}:=A$, $B_{1}=I_{N_{2}}$, and $A_{2}=I_{N_{1}}$. Among these techniques, they discuss a low-rank projection-based scheme, in which they propose constructing extended Krylov subspaces using $A$, following the approach of [33], and employing a standard Krylov basis for the remaining operators. In [27], a low-rank projection-based technique is employed for systems where $N_{2}\ll N_{1}$. The authors opt for the rational Krylov as an approximation space for the Galerkin projection of the original equation, and solve the resulting small-sized equation with either matrix-based Conjugate Gradient or a Kronecker formulation, given that $N_{2}$ is relatively small. In [15], a Krylov-based projection technique is developed for matrix equations of the form (2), where the Krylov basis construction relies on a low-rank decomposition of the operators. The resulting small generalized Sylvester matrix system is solved using the Sherman-Morrison-Woodbury formula through a series of Sylvester equation solves, which leverages the operator’s low-rank structure. Finally, of particular relevance to our study, a basis-update-and-Galerkin (BUG) DLR integrator is proposed in [23] as a low-rank GMRES preconditioner for the generalized Sylvester equation arising from advection-diffusion time-dependent problems with variable coefficients. The approach evolves the basis and the coefficient matrix in the preconditioner using BUG to provide guesses for the basis vectors and coefficient updates. The advantage of the approach is that the approximations involved in the BUG integrator are only present in the preconditioning stage, and not in the converged solution.
In contrast to existing efforts, in this work we propose an optimal implicit time integrator based on SAT adaptive-rank integrators for the variable-coefficient advection-diffusion equation. Instead of evolving the bases and coefficients simultaneously, we evolve them separately to produce an effective, inexpensive implicit solver. In particular, we iteratively construct a set of extended Krylov subspaces (for each dimension) that effectively capture the solution’s range evolution in an iterative manner. A Galerkin projection is then applied to project the full system of equations into a reduced system. The reduced system is solved using Krylov methods like GMRES, with preconditioning based on an approximated operator that transforms the generalized Sylvester equation into a regular Sylvester one, for which suitable efficient solvers exist. To assess the goodness of the solution efficiently, we propose a residual-checking mechanism that leverages the low-rank structure of solutions directly. Such procedure avoids the need to form the solution matrix for norm evaluation. The overall scheme scales as $\mathcal{O}(Nr^{2}+r^{3})$, with $r$ the rank. The main innovations of our proposed algorithm are as follows:
  • Linear “1D-like” Computational Complexity: We introduce a Krylov-based adaptive-rank algorithm for solving equations of the form (2), which arise from implicit integrators for advection-diffusion-type equations. Our method achieves linear computational complexity with respect to $N$ and quadratic complexity with respect to $r$, offering a significant advantage over existing state-of-the-art adaptive-rank approaches. To the best of our knowledge, this is the first study that clearly demonstrates the optimal complexity of $\mathcal{O}(Nr^{2})$ for the advection-diffusion-type equation with variable coefficients.
  • Generalized dimension-wise extended Krylov Subspaces: We propose an averaged-coefficient approximation operator $\tilde{\mathscr{L}}$, which is instrumental not only in building dimension-wise extended Krylov subspaces, but also for preconditioning the reduced projected system. The dimension-wise Krylov subspaces are generated from multiple operators present in ${\mathscr{L}}$, $\tilde{\mathscr{L}}$ and their inverses. Such strategy is numerically demonstrated in this study to be computationally effective with rapid convergence in Krylov iterations.
  • Efficient Preconditioning Strategy: We deploy a GMRES solver for the reduced generalized Sylvester equation from Galerkin projection (of size $r^{2}$), with an Averaging-Coefficient Sylvester (ACS) preconditioner to accelerate the Krylov convergence. The proposed preconditioner is based on a Sylvester direct solver for the averaged-coefficient approximation operator $\tilde{\mathscr{L}}$, which results in a computational complexity of $\mathcal{O}(r^{3})$ (vs. the expected $\mathcal{O}(r^{6})$ of a direct inversion of the Kronecker-product form of the reduced generalized Sylvester equation). This preconditioner is inspired by the recent work in [25], which employed a similar approach for the full-matrix form of the generalized Sylvester equation. A significant enhancement of the algorithm’s efficiency is demonstrated in our numerical tests.
  • Optimized solution representation and low rank residual computation: We employ an SVD-truncated QR procedure for basis augmentation, and an SVD-truncated procedure for the intermediate solution during the Krylov iteration as often as possible. This optimizes the computational efficiency, while preserving the accuracy and structure of the solution. In addition, an efficient low-rank computation of the residual is enabled with $\mathcal{O}(r^{3})$ complexity without explicitly forming the full solution on tensor-product grids. This avoids the computational cost associated with full matrix, which could otherwise spoil the linear computational complexity sought.
@@
The structure of the rest of this paper is as follows. In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the generalized Sylvester equation that results from a finite-difference discretization in space and backward Euler in time. In Section 3, we present our adaptive-rank solution strategy, including the coefficient-averaged approximation operator, construction of extended-Krylov subspaces, preconditioning technique in solving the reduced system, and low-rank computation of the residual norm. The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C. In Section 4, numerical results are presented to demonstrate the efficacy in high-order temporal convergence, adaptive rank evolution of Krylov subspaces, optimal computational complexity, as well as accelerated Krylov convergence. Finally, we conclude with a brief highlight of our proposed algorithm and outlook for future research in Section 5.
@@ 2 Problem Description and Generalized Sylvester Equation
In what follows, we refer to equations in the form of Equation (1) as the Sylvester equation, where the unknown solution matrices are operated on one dimension at a time. Additionally, we use the term generalized Sylvester equation to describe equations in the form of Equation (2), where each unknown solution matrix is simultaneously operated on across both dimensions.
We establish the following notation conventions based on the nature of the symbols. Lowercase Greek letters, such as $\phi(x,y)$ or $\sigma(x,y)$, denote continuous scalar fields. Uppercase bold italic Greek letters, such as $\bm{\mathit{\Phi}}(x,y)$ or $\bm{\mathit{\Sigma}}(x,y)$, are used for vector-valued and tensor-valued functions. Uppercase Greek letters, like $\Phi$ or $\Sigma$, represent discrete matrices arising from the discretization of continuous scalar functions or differential operators on a grid. Boldface uppercase letters, such as $\mathbf{F}$ or $\mathbf{S}$, represent unknown matrix solutions or matrix right-hand sides in matrix equations. The identity matrix of size $n$ is denoted by ${I}_{n}$.
We aim to solve the advection-diffusion equation:
$\frac{\partial f}{\partial t}=\nabla_{v}\cdot(\bm{\mathit{\Phi}}\cdot\nabla_{v}f)-\nabla_{v}\cdot(\bm{\mathit{\Sigma}}f),$ (3)
where $\bm{\mathit{\Phi}}\in\mathbb{R}^{2\times 2}$ is an isotropic diffusion tensor:
@@
and the advection vector $\bm{\mathit{\Sigma}}\in\mathbb{R}^{2}$ is given by:
@@
We assume that the diffusion and advection coefficients are separable and they exhibit a low-rank structure. Specifically, the diagonal diffusion coefficients, $\phi^{x}(x,y)$ and $\phi^{y}(x,y)$, are given by:
$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\quad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$
@@
$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\quad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$
@@
where, in each term, the component independent of the differentiation variable has been pulled out of the differentiation operator. Discretizing this equation using finite differences in space and an implicit method in time naturally leads to a multi-term matrix differential equation where the unknown solution is acted upon from both directions in each term and at the same time. Numerous strategies have been proposed in the literature to solve such equations, most notably the recent work by Palitta and Simoncini in [25]. In this work, inspired by their preconditioning technique [25], we further exploit the low-rank structure of the solution and design a projection-based algorithm with an optimal scaling of $\mathcal{O}(Nr^{2})$, i.e. a linear scaling with respect to the resolution $N$ per spatial dimension, and quadratic scaling with respect to the rank $r$.
2.1 Generalized Sylvester Equation from Implicit Discretization of Advection-Diffusion
To simplify the presentation, we assume rank-one advection and diffusion coefficients, i.e.:
$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\quad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y)^{\top},$
@@ 2.1 Generalized Sylvester Equation from Implicit Discretization of Advection-Diffusion
Thus, Equation (4) simplifies to:
$\frac{\partial f}{\partial t}$ $=\underbrace{\frac{\partial}{\partial x}\left(\phi^{1,x}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}(y)+\phi^{1,y}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}(y)\frac{\partial f(x,y)}{\partial y}\right)}_{\text{diffusion}}$ (5)
${-}\underbrace{\frac{\partial}{\partial x}\left(\sigma^{1,x}(x)f(x,y)\right)\sigma^{2,x}(y){-}\sigma^{1,y}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}(y)f(x,y)\right)}_{\text{advection}}.$
We discretize this equation in a two-dimensional tensor-product spatial grid with ${N}_{1}$ and ${N}_{2}$ grid points in the $x$- and $y$- directions, respectively. The grid points in the $x$- direction are denoted as $x_{i}$ for $i=0,\dots,N_{1}-1$, and similarly for the $y$-direction. We consider the point-wise multiplication of the advection and diffusion coefficients in the discrete by setting:
@@
We approximate the spatial derivative with a three-point stencil, and obtain the following matrix differential equation for $\mathbf{F}$, for which its $(i,j)$ entry $F_{i,j}$ corresponds to the approximation of $f(x_{i},y_{j})$:
$\frac{\partial\mathbf{F}}{\partial t}=\underbrace{D_{xx}^{\Phi}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}{D_{yy}^{\Phi}}^{\top}}_{\text{\emph{diffusion}}}-\underbrace{D_{x}^{\Sigma}\mathbf{F}\Sigma_{2}^{\top}-\Sigma_{1}\mathbf{F}{D_{y}^{\Sigma}}^{\top}}_{\text{\emph{advection}}}.$ (6)
Here, $D_{xx}^{\Phi},D_{yy}^{\Phi},D_{x}^{\Sigma},D_{y}^{\Sigma}$ represent compositions of difference operators with pointwise multiplication operators. See Appendices A and B for more details. For notational simplicity, let
@@
We consider first a backward Euler implicit treatment of the temporal component of the matrix differential equation (8), which gives the following generalized Sylvester equation:
@@
with the residual equation defined as:
$\mathbf{R_{\mathscr{L}}}=\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}=\mathbf{0}.$ (11)
An iterative solver naturally updates the solution $\mathbf{F}_{1}$ until the residual is sufficiently small under a suitable norm.
If we vectorize the right-hand side using the Kronecker product identity, $\text{vec}(A\mathbf{F}B^{\top})=(B\otimes A)\text{vec}(\mathbf{F})$, the second term with a factor of $\Delta t$ in (10) can be written as a matrix-vector multiplication of the form:
$\mathbf{F}\mapsto\left[\underbrace{\left(\Phi_{2}\otimes T_{1}\right)+\left(T_{2}\otimes\Phi_{1}\right)+\left(\Sigma_{2}\otimes T_{3}\right)+\left(T_{4}\otimes\Sigma_{1}\right)}_{\in\mathbb{R}^{N_{1}N_{2}\times N_{1}N_{2}}}\right]\text{vec}(\mathbf{F}),$ (12)
The Kronecker form yields a linear system $Ax=b$, which can be solved either directly or iteratively. The Kronecker approach offers valuable theoretical insights into the relationship between the Kronecker structure and the constructed Krylov basis, a connection exploited in the context of Sylvester equations in [20, 23]. We note that, based on the discussion in [26] on the advantages of the dimension-wise Krylov subspaces projection approach vs. the Kronecker formulation, we choose to adopt the former approach for better computational efficiency and iterative convergence.

### v2

# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients

@@ Abstract
We consider the adaptive-rank integration of 2D and 3D time-dependent advection-diffusion partial differential equations (PDEs) with variable coefficients. We employ a standard finite-difference method for spatial discretization coupled with diagonally implicit Runge-Kutta temporal schemes. The discrete equation is a generalized Sylvester equation (GSE), which we solve with an adaptive-rank algorithm structured around three key strategies: (i) constructing dimension-wise subspaces based on an extended Krylov strategy, (ii) developing an effective preconditioner for the reduced coefficient matrix, and (iii) efficiently computing the residual of the equation without explicitly reverting to the full-rank form. The low-rank decomposition is performed in 2D using SVD, and with high-order SVD (HOSVD) in 3D to represent the tensor in a compressed Tucker format. The computational complexity of the proposed approach is demonstrated numerically to be comparable to the constant-coefficient case [14], scaling as $\mathcal{O}(N{r^{2}}+{r^{d+1}})$ for $d$-dimensional problems (here, $d=2$ or $3$), with $N$ the resolution in one dimension and $r$ the maximal rank during the Krylov iteration (which we find to be largely independent of $N$). We present numerical examples that illustrate the computational efficacy and complexity of our algorithm.
@@
Keywords: Adaptive-rank, extended Krylov subspaces, implicit Runge-Kutta integrators, advection-diffusion equations, Sylvester preconditioning, matrix equations.
1 Introduction:
We propose an adaptive-rank numerical scheme for approximating solutions of time-dependent variable-coefficient advection-diffusion equations. The goal is to reduce the computational complexity of high-dimensional PDE solvers by exploiting the low-rank structure of the solution, building upon existing literature on low-rank PDE solvers [16, 13] and model order reduction (MOR) for solving high-dimensional linear and nonlinear systems [24, 16, 5, 40].
@@ 1 Introduction:
Two primary strategies in the literature exploit evolving low-rank structures for time-dependent PDEs: the dynamical low-rank (DLR) approximation [21, 27, 8, 11, 26] and the step-and-truncate (SAT) method, which has both explicit [23, 10, 18, 17, 19] and implicit [35, 30, 14, 28, 29] variants. Our study employs SAT, discretizing implicitly in time with high-order accuracy, leading to matrix/tensor differential equations to be solved at every time step. For constant coefficients in 2D, the advection-diffusion PDE becomes a Sylvester equation (SE) of the form:
@@
where $\mathbf{X}_{1}\in\mathbb{R}^{N_{1}\times N_{2}}$ represents the solution at time $t_{1}$, $\mathbf{X}_{0}\in\mathbb{R}^{N_{1}\times N_{2}}$ is the initial condition at $t_{0}$, and $A$ and $B$ are difference operators arising from the PDE discretization. Here, $N_{1}$ and $N_{2}$ are the 2D grid dimensions.
Krylov subspace methods have proven effective for solving SEs resulting from the discretization of PDEs [24, 39, 35, 31, 6]. In these methods, the solution bases are chosen from a Krylov (or Krylov-like) subspace of $A$ and $B$ in (1), and the original system is Galerkin-projected onto this reduced subspace. To accelerate the convergence of iterative methods, extensions of Krylov subspaces have been developed [12] involving $A^{-1}$, $B^{-1}$, which have been successfully applied to matrix equations [39, 14].
While substantial progress has been made for constant-coefficient problems, PDEs with variable coefficients remain challenging. Discretizing such systems leads to the so-called generalized Sylvester equation (GSE), of the form:
@@
where $A_{i}$ and $B_{i}$ represent discretization operators for advection and/or diffusion terms. Several methods have been proposed to solve GSEs either with stationary iterative techniques  [5, 38] or Krylov methods  [5, 39, 32, 42]. Stationary iterative methods leverage low-rank techniques to enhance efficiency, achieving computational complexity of $\mathcal{O}(N)$ when implemented effectively [38]. However, their effectiveness critically depends on whether a dominating operator can be correctly identified to enable a contractive iterative process. For the Krylov approach, recent approaches leverage a Galerkin projection of the matrix equation on the dimension-wise Krylov bases when the right-hand side is low-rank [36, 40]. Compared to a standard Kronecker Krylov iteration, dimension-wise Krylov subspaces are inherently richer and yield superior approximations [32, 42, 24].
Several studies in the literature have considered using atypical extended Krylov subspaces to solve GSEs (see [5, 33, 20] and references therein). In [5], the authors present a series of techniques to tackle the Generalized Lyapunov equation, which is a special case of (2) of the form:
$A\mathbf{X}+\mathbf{X}A^{T}+\sum_{j=1}^{m}M_{j}\mathbf{X}M_{j}^{T}=\mathbf{F}.$ (3)
Among them, they discuss a low-rank projection-based scheme, in which extended Krylov subspaces are constructed using $A$ and standard Krylov subspaces are used for the remaining operators. In [20], a Krylov-based projection technique is developed for Generalized Lyapunov equations with $m=1$. The method builds its approximation subspaces by block augmentations generated from $A$ and $M_{1}$. In [33], a low-rank projection-based technique is employed for systems of disparate dimension sizes, $N_{2}\ll N_{1}$. The authors employ a nested sequence of block rational Krylov subspaces generated from all the operators in the $N_{1}$ direction, and solve the resulting small-sized equation with either matrix-based Conjugate Gradient or a Kronecker formulation (tractable since $N_{2}$ is small). In all of these studies, atypical approximation subspaces are formed by augmenting the basis with multiple applications of the operators in the multi‑term matrix equation, a strategy we adopt in this work.
Another Krylov-based line of inquiry, particularly relevant to our study, employs a basis-update-and-Galerkin (BUG) DLR integrator as a low-rank GMRES preconditioner for the GSE arising from anisotropic diffusion time-dependent problems with variable coefficients [28]. The BUG preconditioner evolves the basis and the coefficient matrix to provide guesses for the basis vectors and coefficient updates of the outer GMRES iteration, demonstrating significant potential in controlling GMRES iteration count and maximal Krylov rank. The approach is compatible with high-order temporal integrators. However, the reference offers no discussion of asymptotic computational complexity.
In this study, we propose to solve the GSE for time-dependent advection-diffusion equations leveraging adaptive-rank approximations of solutions on tensor-product grids with $N$ grid points per dimension, and using extended Krylov subspaces to update the basis vectors as well as the rank (denoted as $r$). Our main goal is to minimize the computational cost of the adaptive rank implementation without sacrificing accuracy. Key innovations of our proposed algorithm include:
  • Generalized Extended Krylov (xKrylov) Adaptive-Rank Integrators with Complexity Reduction for 2D and 3D Problems. We consider matrix solutions with factorizations of the form $X=USV^{\top}$, where basis vectors in $U$ and $V$ span the solution space and determine the rank of the solution. In the 2D setting, we employ a truncated singular value decomposition (SVD) to factorize the solution. In 3D, we use high-order singular value decomposition (HOSVD) to represent the 3D tensor in a compressed Tucker format. Our proposed approach constructs a set of xKrylov subspaces (for each dimension) that effectively capture the evolution of solution basis in an iterative manner. A Galerkin projection is then applied to project the full system to a reduced GSE. We propose a residual-checking mechanism that leverages the low-rank structure of solutions directly, thus avoiding forming the full residual matrix for norm evaluations. For the rest of the paper, we refer to the xKrylov iteration as the outer iteration.
  • Effective Preconditioning Strategy: The effectiveness of these projection-based approaches depends on the availability of an effective solver for the reduced GSE (2) of size up to $r$, the maximum size of dimension-wise xKrylov subspaces. There are two main challenges to solve it: (i) an increased computational load due to the multi-term nature of the equation, leading to rapid increase of xKrylov subspace sizes, and (ii) slow convergence of iterative methods (e.g., GMRES) resulting from algebraic ill-conditioning when solved iteratively. These challenges highlight the need for an effective preconditioner. Inspired by [31], we propose an averaged-coefficient Sylvester (ACS) approximation operator to the GSE as a preconditioner to GMRES. The approximate operator is of the Sylvester form and compact-equivalent to the original GSE, ensuring excellent performance as a GMRES accelerator [1, 2]. The direct inversion of the SE features a favorable computational complexity of $\mathcal{O}(r^{d+1})$, to be contrasted with the $\mathcal{O}(r^{3d})$ for $d$-dimensional problems when directly inverting the Kronecker-product form of the GSE. We refer to the preconditioned-GMRES reduced-GSE solve as the inner iteration.
  • Efficient Computational Complexity: As proposed, our method achieves a computational complexity of $\mathcal{O}(Nr^{2}+r^{d+1})$ for $d$-dimensional problems, offering a significant advantage over existing state-of-the-art adaptive-rank approaches. The $Nr^{2}$ term corresponds to the outer xKrylov iteration, and the $r^{d+1}$ term to the inner GMRES iteration, assuming an $r$-independent GMRES convergence rate (which we achieve with the ACS preconditioner). In principle, the rank $r$ may scale with $N$ [24], but we do not find this in practice. The favorable scaling properties of our approach will be verified in our numerical experiments.
@@
The rest of this paper is organized as follows. In Section 2, we consider the case of rank-one advection and diffusion coefficients and derive the GSE that results from a finite-difference discretization in space and backward Euler in time. In Section 3, we introduce an adaptive-rank solution strategy for 2D problems, including the coefficient-averaged approximation operator, construction of xKrylov subspaces, preconditioning technique in solving the reduced system, and low-rank computation of the residual norm. The extension of the technique to high-order temporal discretizations with multi-rank coefficients is provided in Appendix C, and for 3D problems in Appendix D. In Section 4, numerical results are presented to demonstrate the efficacy of the proposed algorithms. Finally, we conclude and provide outlook for future research in Section 5.
@@ 2 Problem Description and Generalized Sylvester Equation
For simplicity, we focus this section on 2D, with the extension to 3D elaborated in Appendix D. We establish the following notation conventions. Lowercase Greek letters, such as $\phi(x,y)$ or $\sigma(x,y)$, denote continuous scalar fields. Uppercase bold italic Greek letters, such as $\bm{\mathit{\Sigma}}(x,y)$ or $\bm{\mathit{\Phi}}(x,y)$, are used for vector-valued and tensor-valued functions. Uppercase Greek letters, like $\Phi$ or $\Sigma$, represent discrete matrices arising from the discretization of continuous scalar functions or differential operators on a grid. Boldface uppercase letters, such as $\mathbf{F}$ or $\mathbf{S}$, represent unknown matrix solutions or matrix right-hand sides in matrix equations. The identity matrix of size $n$ is denoted by ${I}_{n}$.
We aim to solve the time-dependent advection-diffusion equation:
$\frac{\partial f}{\partial t}=\nabla_{v}\cdot(\bm{\mathit{\Phi}}\cdot\nabla f)-\nabla\cdot(\bm{\mathit{\Sigma}}f),$ (4)
where $(x,y)\in\Omega=[a_{1},b_{1}]\times[a_{2},b_{2}]$ and $\bm{\mathit{\Phi}}\in\mathbb{R}^{2\times 2}$ is an isotropic diffusion tensor:
@@
and $\bm{\mathit{\Sigma}}\in\mathbb{R}^{2}$ is the advection vector, given by:
@@
We assume that the diffusion and advection coefficients are separable with finite rank. Specifically, the diagonal diffusion coefficients, $\phi^{x}(x,y)$ and $\phi^{y}(x,y)$, are given by:
$\phi^{x}(x,y)=\sum_{i=1}^{\ell_{x}}\phi^{1,x}_{i}(x)\phi^{2,x}_{i}(y),\quad\phi^{y}(x,y)=\sum_{i=1}^{\ell_{y}}\phi^{1,y}_{i}(x)\phi^{2,y}_{i}(y).$ (5)
@@
$\sigma^{x}(x,y)=\sum_{i=1}^{k_{x}}\sigma^{1,x}_{i}(x)\sigma^{2,x}_{i}(y),\quad\sigma^{y}(x,y)=\sum_{i=1}^{k_{y}}\sigma^{1,y}_{i}(x)\sigma^{2,y}_{i}(y).$ (6)
@@
where, in each term, the factor independent of the differentiation variable has been pulled out of the derivative. To simplify the presentation, we assume for now rank-one advection and diffusion coefficients:
$\phi^{x}=\phi^{1,x}(x)\phi^{2,x}(y),\quad\phi^{y}=\phi^{1,y}(x)\phi^{2,y}(y),$
@@
This will be generalized to multiple ranks below. Equation (7) then simplifies to:
$\frac{\partial f}{\partial t}$ $={\frac{\partial}{\partial x}\left(\phi^{1,x}(x)\frac{\partial f(x,y)}{\partial x}\right)\phi^{2,x}(y)+\phi^{1,y}(x)\frac{\partial}{\partial y}\left(\phi^{2,y}(y)\frac{\partial f(x,y)}{\partial y}\right)}$ (8)
${-}{\frac{\partial}{\partial x}\left(\sigma^{1,x}(x)f(x,y)\right)\sigma^{2,x}(y){-}\sigma^{1,y}(x)\frac{\partial}{\partial y}\left(\sigma^{2,y}(y)f(x,y)\right)}.$
We discretize this equation in a two-dimensional tensor-product spatial grid with ${N}_{1}$ and ${N}_{2}$ grid points in the $x$ and $y$ directions, respectively. The grid points in the $x$ direction are denoted as $x_{i}$ for $i=0,\dots,N_{1}-1$, and similarly for the $y$ direction. We consider the point-wise multiplication of the advection and diffusion coefficients in the discrete by setting:
@@
We approximate the spatial derivative with a second-order finite-difference scheme using a three-point stencil, and obtain the following matrix differential equation for $\mathbf{F}$, for which its $(i,j)$ entry $F_{i,j}$ corresponds to the approximation of $f(x_{i},y_{j})$:
$\frac{\partial\mathbf{F}}{\partial t}={D_{xx}^{\Phi}\mathbf{F}\Phi_{2}^{\top}+\Phi_{1}\mathbf{F}{D_{yy}^{\Phi}}^{\top}}-{D_{x}^{\Sigma}\mathbf{F}\Sigma_{2}^{\top}-\Sigma_{1}\mathbf{F}{D_{y}^{\Sigma}}^{\top}}.$ (9)
Here, $D_{xx}^{\Phi},D_{yy}^{\Phi},D_{x}^{\Sigma},D_{y}^{\Sigma}$ represent compositions of difference operators with pointwise multiplication operators. See Appendices A and B for more details. For notational simplicity, let:
@@
We consider first a backward Euler implicit treatment of the temporal component of the matrix differential equation (11), which gives the following discrete GSE:
@@
and associated residual equation:
$\mathbf{R_{\mathscr{L}}}=\mathbf{F}_{1}-\Delta t\,\mathscr{L}(\mathbf{F}_{1})-\mathbf{F}_{0}.$ (14)
Iterative solvers update the solution $\mathbf{F}_{1}$ until the residual is sufficiently small under a suitable norm. In this work, we opt for Galerkin-projection-based iterative solvers, where the bases spanning $\mathbf{F}_{1}$ are updated by augmenting Krylov-like subspaces, as shown next.